OpenStudy (anonymous):

Generate the first 5 terms of this sequence: (1 point) f(1) = 2 and f(2) = 3, f(n) = f(n - 1) + f(n - 2), for n > 2. 2, 3, 5, 7, 9 2, 3, 4, 7, 11 2, 3, 5, 9, 11 2, 3, 5, 8, 13

OpenStudy (p0sitr0n):

this is the fibonnacci sequence

OpenStudy (p0sitr0n):

it would be 2, 3, 5, 8, 13 2+3=5 3+5=8 5+8=13